The sum of two numbers is 40. One number is 5 more than 4 times the other number. What are the two numbers
Answer by amarjeeth123(569)

Let one number be x.
Then the other number is (40-x).
One number is 5 more than 4 times the other number.
40-x=5+4x
5x=35
x=7
The numbers are 7 and 33.
Verification steps:
33=4(7)+5
So the solution is correct.